OpenStudy (anonymous):

I am going to be a huge fan and I'll see if i can give many medals if you'll help me

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Are you sure cause this is going to take a while

OpenStudy (anonymous):

got all night :)

OpenStudy (anonymous):

alright first are you a guy or girl i say bro a lot

OpenStudy (anonymous):

girl and I do to

OpenStudy (anonymous):

ok lets get this show on the road

OpenStudy (anonymous):

aight bring it on

OpenStudy (anonymous):

For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p?

OpenStudy (anonymous):

it would be f(x)=3p+4

OpenStudy (anonymous):

we would need a input for p

OpenStudy (anonymous):

ther is no imput

OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4).

OpenStudy (anonymous):

well the y would be 4 so assuming there is no input x is 3p so (3p,4)

OpenStudy (anonymous):

and g(4) =11

OpenStudy (anonymous):

inputs are 11 19 16 8

OpenStudy (anonymous):

*answers are

OpenStudy (anonymous):

it is 11 then

OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (anonymous):

answers are 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (anonymous):

I think you may have wrote it wrong because 1 can not be a answer if it is n>2

OpenStudy (anonymous):

lol i didn't write it wrong

OpenStudy (anonymous):

nvm I see my mistake :P

OpenStudy (anonymous):

1,2,3,5,8

OpenStudy (anonymous):

how do you find those

OpenStudy (anonymous):

plug those in for n

OpenStudy (anonymous):

Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(150)?

OpenStudy (anonymous):

what are the statements?

OpenStudy (anonymous):

This is the average number of days the house stayed on the market before being sold for $150,000. The house sold on the market for $150,000 and stayed on the market for an average of 150 days before being sold. The house sold for $150,000. The house stayed on the market for an average of 150 days before being sold.

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

how

OpenStudy (anonymous):

its Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s with replacing the house price instead of p in 1,000s

OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

10

OpenStudy (anonymous):

how

OpenStudy (anonymous):

f(1)+f(2)+f(5-1)

OpenStudy (anonymous):

basicly 2+4+4

OpenStudy (anonymous):

or continue the +2 pattern that seems to be happening

OpenStudy (anonymous):

ya but most teachers like it complicated xD

OpenStudy (anonymous):

If g(x) = 3(x - 2), find the value of x if g(x) = 6.

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

maybe 2.6 x = 2 is not an option

OpenStudy (anonymous):

ya 2.6 just thought they wanted whole numbers

OpenStudy (anonymous):

get some coffee on this little break we could be up till midnight

OpenStudy (anonymous):

is a monster alright

OpenStudy (anonymous):

I only have that

OpenStudy (anonymous):

none of your answrs were right yikes

OpenStudy (anonymous):

hm shame

OpenStudy (anonymous):

ll give you a fan for trying :)

OpenStudy (anonymous):

ha A for effort

OpenStudy (anonymous):

hey what grade are you in

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

hey same :)

OpenStudy (anonymous):

flvs?

OpenStudy (anonymous):

@JDizzle256 you there

OpenStudy (anonymous):

closing it

OpenStudy (anonymous):

aight bye